Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

probability.

Expert replies
by Md.Nazrul Islam » Thu Apr 12, 2012 5:22 am
what is the probability that event E or event F or both will occur .

1)The probability that event E will occur is 0.6

2)The probability that event F will occur is 0.4
Join the discussion
Source: — Problem Solving |

by Shalabh's Quants » Thu Apr 12, 2012 6:36 am
Md.Nazrul Islam wrote:what is the probability that event E or event F or both will occur .

1)The probability that event E will occur is 0.6

2)The probability that event F will occur is 0.4
Given is P(E) = 0.6 & P(F) = 0.4;

=> probability that event E or event F or both will occur will be given by P( E or F or Both)

=> Since P( E or F or Both) = P(E) + P(F) - P(E)* P(F) ;

=> P( E or F or Both)= 0.6 + 0.4 - 0.6*0.4 = 1-0.24 = 0.76.
Shalabh Jain,
e-GMAT Instructor
Join the discussion

by Stuart@KaplanGMAT » Thu Apr 12, 2012 9:32 am
Md.Nazrul Islam wrote:what is the probability that event E or event F or both will occur .

1)The probability that event E will occur is 0.6

2)The probability that event F will occur is 0.4
Hello!

I see that you're fairly new to the forums, so just a friendly tip - please post your questions in the right place! This is the problem solving forum, but this is a data sufficiency question.

Data sufficiency is a test of your knowledge of concepts, NOT your ability to do lots and lots of calculations. Applying common sense and a tiny bit of math knowledge allows us to answer this question quickly and confidently without resorting to any calculations at all.

Step 1 of the Kaplan Method for DS is to analyze the question stem. Too many people just rush into the statements without enough of an understanding of the question itself, often wasting valuable time on unnecessary work.

Q: what's the probability that one or both of two events will occur.

Analysis: we need information about both events - either the probabilities that they will or will not occur.

(1) P(E); nothing about P(F), so insufficient - eliminate A and D.
(2) P(F); nothing about P(E), so insufficient - eliminate B.

Since neither statement was sufficient alone, we have to combine.

Together: P(E) and P(F).
We think: with the two individual probabilities, we can answer ANY question about the two events. SUFFICIENT, choose (C).

Remember: in DS we don't care WHAT the answer is, we only care WHETHER IT'S POSSIBLE to answer the question. Don't waste your time on unnecessary calculations.

* * *

Aside #1: when you're done a DS question, it's useful to ask yourself "if this were a PS question, what would have been the quickest way to solve it?" So, let's go into "review mode" and find the best solution for this problem.

An often time-saving approach to complex probability questions is the "one minus" method. Since the sum of all probabilities is 1, we can form the equation:

Prob(what you want) = 1 - Prob(what you don't want).

In this problem, we get:

Prob(one or both events) = 1 - Prob(neither event)

Since P(E) = 0.6, P(not E) = .4
Since P(F) = 0.4, P(not F) = .6

Accordingly, Prob(neither E nor F) = .4 * .6 = .24,

and

Prob(one or both) = 1 - .24 = .76

* * *

Aside #2: there's a problem with the wording of the question that actually disqualifies it from the GMAT - nowhere does it say that events E and F are independent. We're assuming that they are, but the GMAT would never leave a question open to multiple interpretations. So, the question should have read "what is the probability that one or both of independent events E and F occur?"
Image

Stuart Kovinsky | Kaplan GMAT Faculty | Toronto

Kaplan Exclusive: The Official Test Day Experience | Ready to Take a Free Practice Test? | Kaplan/Beat the GMAT Member Discount
BTG100 for $100 off a full course
Join the discussion